OpenStudy (gabylovesyou):

What is the equation in standard form of the line which passes through (4, - 2) and has a slope of -3?

OpenStudy (gabylovesyou):

@Nurali

OpenStudy (gabylovesyou):

@sleepyjess

OpenStudy (sleepyjess):

I am thinking we should first get point slope form then convert that to standard form.

OpenStudy (sleepyjess):

So point slope form would be y + 2 = -3(x - 4)

OpenStudy (gabylovesyou):

ok... -3 * x = -3x and -3 * 4 = -12 so y + 2 = -3x - 12 ?

OpenStudy (sleepyjess):

Yep, then we need to get y to the right side.

OpenStudy (gabylovesyou):

add 3 ?

OpenStudy (nurali):

We use the formula: y - y1 = m(x - x1) y - (-2) = -3(x - 4) y - (-2) = y + 2 Now we have: y + 2 = -3(x - 4) Cross multiply: (y + 2) = -3(x - 4) Therefore: y + 2 = -3x +12 y=-3x+12-2 y=-3x+10

OpenStudy (sleepyjess):

We could also add 3x. I think that is what you were trying to say?

OpenStudy (sleepyjess):

That is slope-intercept form though...

OpenStudy (gabylovesyou):

wait what

OpenStudy (sleepyjess):

@Nurali did slope-intercept form. They are asking for standard form which is c = ax + by

OpenStudy (gabylovesyou):

oh ok.. so 3x + y + 2 = -12

OpenStudy (sleepyjess):

Yeah, that works :) now just subtract 2 and we're done ^_^

OpenStudy (gabylovesyou):

3x + y = -14

OpenStudy (sleepyjess):

Wait, it should be +12

OpenStudy (gabylovesyou):

o-o ohh

OpenStudy (sleepyjess):

-3*-4 = 12

OpenStudy (nurali):

3x+y=10

OpenStudy (gabylovesyou):

ohhhhhhhhh 3x + y = 10... mkay cx

OpenStudy (gabylovesyou):

thanks

OpenStudy (sleepyjess):

\(\LARGE\cal\color{#ff0074}{No~Problem!}\\\bbox [10pt, #00ffab,border:5pt solid blue ]{\color{#872AF0}{\huge\cal ~\dagger~\heartsuit sleepyjess~\heartsuit~\dagger}}\)